Which of the following statements is false?
(a) In a balanced Wheatstone bridge, if the cell and the galvanometer are exchanged, the null point is disturbed
(b) A rheostat can be used as a potential divider
(c) Kirchhoff’s second law represents energy conservation
(d) Wheatstone bridge is the most sensitive when all the four resistances are of the same order of magnitude

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

Statement A is false since after exchange of cell and galvanometer in the Wheatstone bridge, the condition that\dfrac{R_1}{R_3}=\dfrac{R_2}{R_4}must be satisfied still holds true since in the eariler configuration\dfrac{R_1}{R_2}was equal to\dfrac{R_3}{R_4}.